# LLM-Agent-UMF: LLM-based Agent Unified Modeling Framework for Seamless Integration of Multi Active/Passive Core-Agents

### v2

@@ 3.2.2 Memory Module
The memory module is responsible for the storage and retrieval of information pertinent to the core-agent’s activities, thereby enhancing the core-agent’s decision-making efficiency and task execution capabilities. In [27] and [5], the memory module in an LLM-based agent was approached from an abstract functional perspective, neglecting its analysis from a software architectural viewpoint. This led to some overlap between the memory module and the other modules defined in the framework suggested by [5]. Consequently, we propose a more comprehensive and well-defined presentation of this module based on three perspectives: Memory Scope, Memory Location, and Memory Format (Figure 5).
@@
The framework presented in the survey [5] classified the memory structure into Unified Memory, intended to emulate human short-term memory via in-context learning, and Hybrid Memory, which represents both short-term and long-term memory functionalities. While these names aimed to differentiate types of agent systems, they deviate unnecessarily from the well-established terminology in the field. A more conventional and widely recognized categorization is introduced through the first perspective, Memory Scope, which includes short-term and long-term memory, aligning more with the human memory types as detailed in [28].
As opposite to how it was defined in framework [5], our short-term memory definition diverges to focus more on the impact of the core-agent rather than on the LLM. From a content perspective, various sources of information contribute to short-term memory. As pointed out in paper [27], data can be derived from one trial of a given task or from previous trials of the same task, which is referred to as short-term memory. This data has a narrow scope that focuses on a specific task and primarily communicated to the LLM for the purpose of in-context learning and enhancing the LLM capability with more information related to the task at hand. Conversely, long-term memory refers to the ability to store and recall information over extended periods, beyond the scope of a specific task. This type of memory enables the core-agent to maintain coherence and context over prolonged interactions, learning and adapting from past experiences. Namely, MemoryBank [29] stores all interactions with user in a large symbolic memory and processes experiences into high-level summaries to reflect upon future similar tasks.
However, there is always a limit to the amount of memory a core-agent can retain. Therefore, techniques such as forgetting mechanisms must be implemented to decide which memories to discard and which to retain [27]. As mentioned earlier, survey [5] treats memory from an abstract functional aspect, which is inconsistent with our framework that emphasizes a clear delineation of distinct modules. Our approach aims to provide a more precise and structured understanding of core-agent components from a software perspective. For instance, according to survey [5], the memory module includes a "reflection" operation with a cognitive aspect that we believe belongs to the planning module. The planning module focuses on achieving the agent’s goals and is responsible for decision-making and synchronizing among all other modules. Thus, it makes it more suitable for handling reflections and memory optimization by collaborating with the memory module to access stored data. Likewise, extracting useful information from memory and reflecting upon it to produce a solid plan belongs to planning module responsibilities. For example, Voyager [30] considers environmental feedback, handled as short-term information, and leverages the LLM capability to adjust its plan and make more efficient and rational decisions in the scope of planning module.
Furthermore, the writing operation is an essential aspect of a memory, distinguishing it from data repositories. Therefore, we exclude knowledge repositories from the memory category. As a matter of fact, the retrieval of knowledge from databanks, in a read-only way, falls under the responsibilities of the action module, which will be discussed in Section 3.2.4. As we have set writing and reading as essential operations for the memory module and excluded reflection from the possible set of operations, we decided to omit the Memory Operation perspective discussed in survey [5] from our framework.
Beyond focusing on the scope of the data, inside trial and across trials, as stated in [27], we emphasize the location of the memory because it is more relevant from a software architecture point of view. Hence, we introduce the second perspective, Memory Location, which encompass two categories: Embedded Memory, internal to the core-agent boundaries, and Memory Extension, outside the core-agent boundaries, yet still within the boundaries of the agent system as depicted in Figure 3.
Finally, the last perspective, Memory Format, already present in both works [5] and [27], focuses on the shape and the representation of the memory that could have diverse manifestations: natural language, embeddings, SQL databases, or structured lists. In fact, memory information such as agent behaviors and observations could be directly described using raw natural language, providing flexibility and retaining rich semantic information. Another solution would be to encode memory information into embedding vectors to enhance retrieval and reading efficiency. In addition, as exemplified by ChatDB [31], databases could be used as memory holders to store memory information in a structured representation, allowing agents to manipulate memories efficiently and comprehensively using SQL queries. And finally, structured lists, as used in GITM [32] let core-agents save the sequential actions of subgoals in a structured way and organize memory information into hierarchical lists to convey semantic information concisely.
As an observation, the most common format is textual, nevertheless [5] notes that these formats are not mutually exclusive, so one core-agent can handle multiple formats, such as GITM’s key-value list structure that combines embedding vectors and raw natural language, to harness the respective benefits of each approach.
@@ 3.2.3 Profile Module
Similarly to the approach presented by the paper [5], our framework defines the function of the profile module as to establish the role of the LLM and adopts more diverse methods. This template explicitly separates the roles of both the core-agent and the LLM. Indeed, the profile module facilitates the dynamic adaptation of various profiles tailored to specific use cases and strategies employed by the planning module. The Profile module features four methods for defining profiles: the Handcrafted In-Context Learning Method, the LLM-generation method, the Dataset Alignment Method, and the newly introduced Fine-tuned Pluggable Modules method (Figure 6).
@@
The Handcrafted In-Context Learning Method, previously referred to as the Handcrafting Method in the survey [5], involves deputing the core-agent to set the LLMs profiles through in-context learning techniques that employ pre-configured prompts. This method allows for fine-grained control over the LLM’s personality and behavior. While it is a straightforward method to implement, it necessitates the use of LLMs that are well-suited for in-context learning and often possess a cumbersome size.
The LLM-Generation Method facilitates the automatic creation of profiles for agents using LLMs. This method begins by specifying the profile’s characteristics which include detailed information such as age, gender, and interests. Optionally, several seed agent profiles can be selected to serve as few-shot examples, as outlined in the paper [5]. Once these seed profiles are established, LLMs are employed to generate additional profiles by referring to the initial seed examples. For instance, RecAgent [33] suggests designing appropriate prompts that encourage a GPT model to generate comprehensive profile descriptions by referring to a table of attributes corresponding to various samples and generate additional profiles. The LLM-Generation Method, while offering significant time-saving advantages, is constrained by its reliance on LLMs. This dependence can lead to potential biases [34] or inaccuracies in generated agent profiles.
Additionally, the Dataset Alignment Method derives profiles from real-world datasets [5], which consists of data about actual individuals. This approach starts by organizing the information in these datasets into natural language prompts that describe the characteristics of the role of the LLM. These structured data are then used to create the profiles. In the study conducted by [35], researchers utilized GPT-3 alongside real world demographic data from ANES to assign roles based on characteristics such as state of residence. Subsequently, they evaluated whether GPT-3 could mimic real human behavior reliably. The Dataset Alignment Method ensures that the LLM profile accurately reflects real-world attributes and behaviors, thereby making it meaningful and realistic. However, the effectiveness of this method relies heavily on the accuracy and representativeness of the underlying real-world datasets.
Finally, we introduced the Fine-tuned Pluggable Modules Method, a pioneering solution to set LLM’s profile leveraging several state-of-the-art techniques, aiming to provide an efficient customization and adaptation of the LLM’s profile. This approach defines the LLM profile by injecting a pluggable module fine-tuned to influence the language model behavior [36]. Indeed, that module is a set of additional tunable parameters that must be previously trained using Parameter-Efficient Fine-Tuning (PEFT) techniques, such as Sequential adapter (AdapterS), Prompt-tuning, or LoRA [37].
This process ensures precise and effective customization to achieve the desired profile by eliminating the need for in-context learning, which traditionally involves injecting extensive prompt information into each query. By omitting this step, the required context size is significantly reduced, as the adapters directly encode the desired behaviors and profiles into the model parameters. This reduction in context size and memory footprint allows the LLM to operate more efficiently, using less memory during inference without compromising performance and flexibility.
@@ 3.2.4 Action Module
While we have enriched the approaches applied in this module, we define it within the LLM-Agent-UMF almost similarly to the methodology described in [5]. The action module is responsible for converting high-level instructions from the planning module into low-level actions, leveraging tools available in its environment. In our framework, the action module interacts with the security module to ensure that any executed action aligns with predefined criteria and prevent information leakage as will be explained further in Section 3.2.5. In essence, our analysis considers four fundamental aspects: Goal, Trigger, Action Space, and Impact (Figure 7).
@@
The first perspective, the Action Goal, represents what the core-agent intends to accomplish by performing actions based on various objectives such as Task Completion, Communication, or Environment Exploration which aligns with the framework proposed in [5]. The second aspect, Action Trigger, focuses on how the actions are produced and the catalysts behind them. This perspective is derived from the Action Production perspective originally proposed in [5], but we emphasize more on the causality that drives action production. In this regard, the "Action via Memory Recollection" approach mentioned in [5] is considered to be better handled by the planning module as it focuses on leveraging past experiences for making appropriate decisions. In fact, the core-agent can produce actions as a result of two primary triggers: Plan Following and API Call Request, newly introduced in our framework. In the case of Plan Following trigger, the core-agent can execute actions by adhering to pre-generated plans elaborated by the planning module. As a matter of fact, in GITM [32], the agent breaks down the task into subtasks and procedures and executes the appropriate actions in the Minecraft world via the LLM interface. By exploring other state-of-the-art techniques like TALM [6] and ToolFormer [38] that exemplify LLMs’ capacity to improve performance across diverse tasks through incorporating external tools, we introduce a new trigger named API Call Request. This approach enables the core-agent to execute actions in response to API call requests initiated by the LLM, facilitating smooth integration and effective utilization of external resources.
@@
The third perspective, Action Space, defines the set of possible actions that can be performed by the core-agent in response to its objectives and environmental factors. This concept is distinct from operations involving internal state changes and memory management, which are handled separately by the Memory Module. In contrast to [5], our framework emphasizes this separation to avoid potential overlap between action and memory modules functions, thus ensuring a clear division of responsibilities for efficient execution. The action module can leverage tools ranging from basic software components like calculators to more advanced systems accessed through API calls, as exemplified in the HuggingGPT work that leverages the HuggingFace API to accomplish complex user tasks [39]. Alternatively, the core-agent may expand its operational scope and knowledge by communicating with read-only data sources such as knowledge repositories. This process is often referred to as Retrieval-Augmented Generation (RAG) [40]. This is particularly crucial for AI systems deployed in critical industries like healthcare where explainability in decision-making processes becomes paramount [41, 42]. By interacting with these external resources, transparency, provenance and traceability are ensured to guarantee the reliability of the agent’s outputs [43]. It is useful to note here that data repositories can have the same formats discussed in the memory module (Section 3.2.2). For instance, LLamaIndex [44] stores data as vector embeddings at the indexing stage to leverage semantic search, where the similarity between embeddings is used to rank documents by their relevance to a query. In contrast, ReAct [45] uses a textual data repository, like Wikipedia, to mitigate error propagation in chain-of-thought reasoning.
Finally, Action Impact refers to the consequences resulting from an action. Numerous impacts can be cited, such as changing the environment by moving to different locations, gathering resources, or synthesizing new chemical compounds [12]. Actions can also result in the alteration of the internal state of the agent, especially when acquiring new knowledge and collaborating with the memory module. Additionally, actions can trigger new ones, creating a chain of actions during task execution.
@@ 3.2.5 Security Module
As LLMs continue to advance and become increasingly prevalent in various applications, it becomes crucial to address potential risks, ethical considerations and unintended consequences associated with their deployment [46]. These concerns revolve around issues such as unauthorized or unethical use, data biases, privacy breaches, and the spread of misinformation [4]. This has led to the introduction of guardrails recently in LLMs field as algorithms to identify and prevent the misuse of LLMs [13]. To bolster our framework for trustworthy AI systems, we propose a fifth module: the Security Module. This module aims to provide a more capable and responsible core-agent. Its role is monitoring the action module specifically in production environments to ensure the safety and responsible use of LLMs.
The Security Module operates within the parameters of the Confidentiality, Integrity, Availability (CIA) triad [47], a crucial model which encompasses three pivotal principles in the security field. Confidentiality is centered around protecting sensitive information from unauthorized access or disclosure. Within LLMs-based agents, this principle is critical in safeguarding user data and ensuring the non-divulgence of sensitive information. Integrity is concerned with the accuracy, consistency, and trustworthiness of data throughout its lifecycle. For agents, this principle involves maintaining the reliability of the model’s outputs and preventing any unauthorized modifications to the system or its data. Lastly, availability focuses on ensuring that information and resources are accessible to authorized users whenever required. In the context of LLM applications, this entails maintaining system uptime and providing safe responses to user inquiries. By adhering to these principles, the Security Module aims to establish a robust and trustworthy environment for the operation of the agent, effectively addressing critical concerns related to their deployment and usage.
@@
While conducting a thorough research around the Security Module, our approach will involve exploring multiple facets: Identifying and securing critical assets and data within the core-agent modeling framework, implementing strategies and mechanisms to protect these assets from potential threats, ensuring the core-agent’s ability to effectively respond to and mitigate any security incidents, and maintaining the privacy of user data while adhering to relevant data protection regulations (Figure 8).
Security measures:
Regardless of the guardrail type deployed, the Security Module encompasses three fundamental axes: Prompt safeguarding, response safeguarding, and data privacy safeguarding.
@@
    Prompt safeguarding necessitates employing measures to detect and mitigate unauthorized access to Large Language Models through prompt injection attacks [48]. Techniques for enhancing the security of LLM-based agents can be integrated directly into the LLM itself, with Adversarial Training (AT) being a prominent example [49] AT enhances an LLM’s defense mechanisms by fine-tuning it with augmented training data containing adversarial examples, thereby increasing the model’s ability to safeguard against malicious prompts and improving its robustness. However, AT faces significant limitations, such as the challenges in efficiently selecting adversarial examples and the model’s exposure to adversarial perturbations such as HOUYI [50], a black-box prompt injection attack. Moreover, such training-based security techniques may impact the generative performance of the LLM which necessitates additional evaluation steps.
    Other techniques address these limitations by decoupling the security measures from the LLM and delegating them to a distinct entity that we identify as a core-agent supplemented with a security module as illustrated in Figure 3. This decoupling is essential for improved protection and enables the implementation of more advanced security protocols on the LLM input, which can evolve independently of the LLM’s training process. An example of this approach is Nvidia NeMo [51], which functions as an intermediary layer between users and LLMs, employing advanced techniques such as vector databases and comparison with stored canonical forms to filter and process user inputs before they reach the model, thereby providing robust prompt safeguarding without directly modifying the underlying LLM.
    These approaches are essential to address scalability challenges, enable proactive defense, and facilitate continuous learning in LLM security, ensuring that protection mechanisms can adapt to new threats and maintain the integrity of LLM interactions.
@@
    The recent survey [52] has demonstrated that despite the implementation of prompt safeguarding techniques, the overall resilience of Large Language Models (LLMs) against advanced attacks, known as jailbreaks, may not experience significant improvement. These jailbreaks are designed to exploit biases or vulnerabilities within language models by manipulating their responses. Notable examples include white-box attacks AutoDAN-Zhu [53], which generate stealthy prompts to avoid triggering the model protective mechanisms. Additionally, black-box attacks leverage manually crafted prompts to deceive the LLM [52]. The effectiveness of these jailbreak techniques is further illustrated in [13], where researchers successfully achieved a jailbreak attack on ChatGPT 3 by framing potentially harmful query, "how to hotwire a car" as a hypothetical scenario. The existence of these jailbreak methods highlights the urgent necessity for continuous and rigorous monitoring of LLM outputs to detect and mitigate potential breaches. It is useful to note that the aforementioned jailbreak attack was addressed and resolved in later versions of ChatGPT, 3.5 and 4, as confirmed by [13].
@@
    However, under our proposed framework, the primary objective shifts from solely protecting the privacy of training data to ensuring that the LLM does not divulge sensitive information to external tools. This approach aims to maintain the privacy and security of data while interacting with other systems. Indeed, as previously mentioned, the core-agent can leverage external tools, APIs, and knowledge repositories to augment the capabilities of the LLM. Being part of the agent as a whole system, internal tools are part of the privacy circle, thereby, there is no need to apply security measures on communication procedures. However, the use of external resources introduces potential risks that require specific attention. These risks include a lack of robust data privacy measures, potentially leading to data leaks or unauthorized access. For example, when the core-agent utilizes third-party services to access additional information or perform specific tasks, the data transmitted may contain sensitive details that specific systems are not authorized to access. Additionally, such sensitive information could be intercepted or mishandled if proper secure channels are not leveraged.
@@
Thus, the security module operates along three fundamental axes: Prompt Safeguarding, Response Safeguarding, and Data Privacy Safeguarding. Their integration forms a robust defense mechanism capable of mitigating diverse threats, ranging from prompt injection attacks to potential data breaches. Through this approach, the security module shapes the behavior of the core-agent, prioritizing security in every aspect, from data retrieval to external communication.
@@
To implement these security axes effectively, various guardrail methodologies have been developed. These guardrails act as the operational layer of the security module, translating high-level security objectives into actionable safeguards. The paper [52] delves into diverse guardrail methodologies and solutions offered by LLM service providers and the open-source community. Through meticulous analysis of these methodologies, two primary types emerge, rule-based guardrails and LLM-based guardrails.
@@
To wrap up the LLM-based agent unified modeling framework, our proposed solution emulates the modular architecture of the human brain by introducing the core-agent component, which encompasses five internal modules: planning, memory, profile, action, and security. This modular design introduces enhancements on the security level and addresses challenges related to extendibility and maintainability, effectively separating the core-agent functionalities from the LLM. Consequently, this structured aspect highlights the role of each module and the implication of their integration, especially the planning module as it gives core-agents adopting it an authoritative aspect. The latter perspective leads to the classification of core-agents that will be discussed in Section 3.3 .

### v3

@@ 4.2.2 Memory Module
The memory module manages the storage and retrieval of information critical to the core-agent’s activities, supporting efficient decision-making and task execution. Unlike read-only knowledge repositories, which are handled by the action module (Section 4.2.4), the memory module supports both reading and writing operations to dynamically update task-related data. It is organized around three perspectives: Memory Scope, Memory Location, and Memory Format (Figure 5).
@@
The Memory Scope perspective categorizes memory into short-term and long-term types, aligned with human memory models [18]. Short-term memory stores data from one or multiple trials of a specific task [22], enabling in-context learning to enhance LLM performance on immediate tasks. Long-term memory retains information over extended periods, maintaining coherence across interactions and adapting from past experiences. For example, MemoryBank [49] stores user interactions in a symbolic memory, summarizing experiences for future tasks. Due to memory constraints, forgetting mechanisms are employed to selectively retain or discard data [22].
The Memory Location perspective distinguishes between Embedded Memory, stored within the core-agent, and Memory Extension, stored externally but within the agent system (Figure 3). This categorization supports flexible memory management tailored to architectural needs.
The Memory Format perspective defines how memory is represented, including natural language, embeddings, SQL databases, or structured lists. Natural language offers semantic richness, while embeddings improve retrieval efficiency. For instance, ChatDB [23] uses SQL databases for structured memory manipulation, and GITM [24] employs structured lists to organize sequential actions hierarchically. Multiple formats can be combined, as in GITM’s key-value lists blending embeddings and natural language, to leverage their respective strengths.
@@ 4.2.3 Profile Module
The profile module, an integral component of the core-agent’s internal structure, defines the role of the Large Language Model (LLM) while clearly distinguishing it from the core-agent’s coordinating responsibilities. This separation ensures that the LLM operates with a specific persona tailored to the task, enhancing the core-agent’s ability to adapt dynamically to diverse use cases and planning strategies. The module employs four distinct methods to create these profiles, as illustrated in Figure 6: Handcrafted In-Context Learning, LLM-Generation, Dataset Alignment, and Fine-tuned Pluggable Modules.
@@
The Handcrafted In-Context Learning Method enables the core-agent to configure LLM profiles using pre-configured prompts, providing fine-grained control over the LLM’s personality and behavior. This approach leverages in-context learning techniques, requiring LLMs that are adept at interpreting such prompts, often necessitating models of substantial size. While straightforward to implement, its reliance on large LLMs can pose computational challenges, making it resource-intensive but effective for precise customization.
The LLM-Generation Method automates profile creation by harnessing the LLM’s generative capabilities. It begins by specifying profile characteristics, such as age, gender, or interests, and may incorporate seed profiles as few-shot examples to guide generation. For example, RecAgent [50] designs prompts that encourage a GPT model to generate comprehensive profile descriptions based on attribute tables, streamlining the process. However, this method’s dependence on LLMs introduces risks of biases [51] or inaccuracies, necessitating careful validation of generated profiles.
The Dataset Alignment Method derives profiles from real-world datasets containing individual data, ensuring realistic and grounded LLM roles. Information from these datasets is organized into natural language prompts that describe the LLM’s role characteristics. In a study by [52], researchers used GPT-3 with ANES demographic data to assign roles based on attributes like state of residence, mimicking real human behavior reliably. This method’s effectiveness relies heavily on the accuracy and representativeness of the datasets, making data quality a critical factor.
The Fine-tuned Pluggable Modules Method, a novel approach, customizes LLM profiles by injecting fine-tuned modules that adjust model behavior [43]. These modules, trained using Parameter-Efficient Fine-Tuning (PEFT) techniques such as Sequential Adapter, Prompt-tuning, or LoRA [53], encode desired profiles directly into model parameters. This eliminates the need for extensive in-context learning prompts, reducing context size and memory footprint while maintaining flexibility and performance, offering an efficient and precise customization solution.
@@ 4.2.4 Action Module
The action module, a core component of the core-agent, translates high-level instructions from the planning module into executable low-level actions by leveraging tools within its environment. It collaborates closely with the security module to ensure that actions adhere to predefined integrity, safety and privacy criteria, as discussed in Section 4.2.5. The module’s operations are structured around four key perspectives: Action Goal, Action Trigger, Action Space, and Action Impact, as depicted in Figure 7.
@@
The Action Goal perspective outlines the objectives the core-agent aims to achieve through its actions. These objectives include Task Completion, such as solving a specific problem; Communication, such as interacting with users or other systems; and Environment Exploration, such as navigating a virtual or physical space. This perspective ensures that each action is purposefully aligned with the core-agent’s intended outcomes, providing a clear direction for the execution process.
The Action Trigger perspective identifies the catalysts that initiate actions, encompassing two primary triggers: Plan Following and API Call Request. Plan Following involves executing actions based on pre-generated plans from the planning module. For instance, in GITM [24], the agent decomposes tasks into subtasks and executes actions in the Minecraft world via an LLM interface. The API Call Request trigger, inspired by techniques like TALM [6] and ToolFormer [38], enables actions in response to LLM-initiated API calls, facilitating seamless integration with external resources.
@@
The Action Space perspective defines the set of possible actions that can be performed by the core-agent in response to its objectives and environmental factors. This concept is distinct from operations involving internal state changes and memory management, which are handled separately by the Memory Module. The action module can leverage tools ranging from basic software components like calculators to more advanced systems accessed through API calls, as exemplified in the HuggingGPT work that leverages the HuggingFace API to accomplish complex user tasks [54]. Alternatively, the core-agent may expand its operational scope and knowledge by communicating with read-only data sources such as knowledge repositories. This process is often referred to as Retrieval-Augmented Generation (RAG) [55]. This is particularly crucial for AI systems deployed in critical industries like healthcare where explainability in decision-making processes becomes paramount [56, 57]. By interacting with these external resources, transparency, provenance, and traceability are ensured to guarantee the reliability of the agent’s outputs [58]. It is useful to note here that data repositories can have the same formats discussed in the memory module (Section 4.2.2). For instance, LlamaIndex [59] stores data as vector embeddings at the indexing stage to leverage semantic search, where the similarity between embeddings is used to rank documents by their relevance to a query. In contrast, ReAct [60] uses a textual data repository, like Wikipedia, to mitigate error propagation in chain-of-thought reasoning.
The Action Impact perspective captures the consequences of actions, which may include altering the environment, such as moving locations, gathering resources, or synthesizing chemical compounds [12]. Actions can also modify the core-agent’s internal state, particularly when acquiring new knowledge. This perspective highlights the tangible and intangible effects of actions, underscoring their role in achieving the core-agent’s goals and adapting to dynamic contexts.
@@ 4.2.5 Security Module
The security module addresses risks in Large Language Model (LLM) deployment, such as unauthorized or unethical use, data biases, privacy breaches, and misinformation [61, 4]. It defines security as specifying what is permitted, by whom, when, and how, serving as the core-agent’s enforcement mechanism, while safety focuses on preventing harm. Aligned with standards like the EU AI Act [25] and ISO/IEC 22989:2022 [26], which emphasize robustness and cybersecurity, the module continuously monitors the action module to enforce security and safety requirements. Incorporating guardrails—specialized algorithms that detect and mitigate misuse [13]—it ensures robust protection for LLM-based systems. The module integrates the Confidentiality, Integrity, and Availability (C.I.A.) triad [27], operationalizing these principles through two key perspectives: Safeguarding Measures (Prompt, Response, and Data Privacy Safeguarding) and Guardrail Types (rule-based and LLM-based guardrails), as illustrated in Figure 8, operating independently of cognitive functions like planning and memory to maintain robust protection.
Security Measures:
@@
The security module implements three safeguarding mechanisms—Prompt Safeguarding, Response Safeguarding, and Data Privacy Safeguarding. These mechanisms, applied at input and output levels, form a layered defense to mitigate threats like prompt injections and data breaches.
@@
    Prompt safeguarding necessitates employing measures to detect and mitigate unauthorized access to Large Language Models (LLMs) through prompt injection attacks [62]. Techniques for enhancing the security of LLM-based agents can be integrated directly into the LLM itself, with Adversarial Training (AT) being a prominent example [63]. AT enhances an LLM’s defense mechanisms by fine-tuning it with augmented training data containing adversarial examples, thereby increasing the model’s ability to safeguard against malicious prompts and improving its robustness. However, AT faces significant limitations, such as challenges in efficiently selecting adversarial examples and vulnerability to adversarial perturbations like HOUYI [64], a black-box prompt injection attack. Moreover, such training-based security techniques may impact the generative performance of the LLM, necessitating additional evaluation steps.
    To address these limitations and enhance protection, other techniques decouple security measures from the LLM, delegating them to a distinct entity identified as a core-agent supplemented with a dedicated security module, as illustrated in Figure 3. This decoupling is critical for robust prompt safeguarding and enables the implementation of advanced security protocols that continuously monitor and filter LLM inputs, evolving independently of the LLM’s training process. An example of this approach is Nvidia NeMo [29], which functions as an intermediary layer between users and LLMs, employing advanced techniques such as vector databases and comparison with stored canonical forms to detect and mitigate malicious inputs before they reach the model, thereby providing robust prompt safeguarding without directly modifying the underlying LLM.
@@
    The recent survey [28] has demonstrated that despite the implementation of prompt safeguarding techniques, the overall resilience of Large Language Models (LLMs) against advanced attacks, known as jailbreaks, may not experience significant improvement. These jailbreaks are designed to exploit biases or vulnerabilities within language models by manipulating their responses. Notable examples include white-box attacks AutoDAN-Zhu [65], which generate stealthy prompts to avoid triggering the model protective mechanisms. Additionally, black-box attacks leverage manually crafted prompts to deceive the LLM [28]. The effectiveness of these jailbreak techniques is further illustrated in [13], where researchers successfully achieved a jailbreak attack on ChatGPT 3 by framing potentially harmful query, “how to hotwire a car” as a hypothetical scenario. The existence of these jailbreak methods highlights the urgent necessity for continuous and rigorous monitoring of LLM outputs to detect and mitigate potential breaches. It is useful to note that the aforementioned jailbreak attack was addressed and resolved in later versions of ChatGPT, 3.5 and 4, as confirmed by [13].
@@
    However, under our proposed framework, the primary objective shifts from protecting the privacy of training data to ensuring that the LLM does not divulge sensitive information to external tools. This approach aims to maintain the privacy and security of data while interacting with other systems. Indeed, as previously mentioned, the core-agent can leverage external tools, APIs, and knowledge repositories to augment the capabilities of the LLM. Being part of the agent as a whole system, internal tools are part of the privacy circle, thereby, there is no need to apply security measures on communication procedures. However, the use of external resources introduces potential risks that require specific attention. These risks include a lack of robust data privacy measures, potentially leading to data leaks or unauthorized access. For example, when the core-agent utilizes third-party services to access additional information or perform specific tasks, the data transmitted may contain sensitive details that specific systems are not authorized to access. Additionally, such sensitive information could be intercepted or mishandled if proper secure channels are not leveraged.
@@
The security module operates across three axes: Prompt Safeguarding, Response Safeguarding, and Data Privacy Safeguarding. These integrated mechanisms mitigate threats like prompt injection and data breaches, providing layered protection at input and output levels. This approach prioritizes security in the core-agent’s operations, from data retrieval to external communication.
@@
To implement the security axes, the security module uses guardrail methodologies as its operational layer, translating security objectives into actionable safeguards. Two primary types are rule-based guardrails and LLM-based guardrails, as outlined in [28].